# Anchor Token Matching: Implicit Structure Locking for Training-free AR Image Editing

### v1

@@ 2 Related Work
Text-guided Image Editing. Text-guided image editing aims to modify image content based on semantic prompts while preserving irrelevant regions. Traditional GAN-based methods [16], including StyleCLIP [35] and CLIPStyler [24], achieve local editing by optimizing in the GAN latent space with CLIP guidance, but their performance is constrained by the capacity of pre-trained GANs. Recent diffusion-based approaches have become mainstream, yet they often require complex inversion and latent optimization processes to balance fidelity and controllability. For example, optimization-based inversion methods [30, 53, 26, 17] refine latent noise for accurate reconstruction and manipulate cross-attention maps (e.g., Prompt-to-Prompt [18]) to preserve structure. InstructPix2Pix [2] bypasses inversion by training on paired edit data, but relies on Prompt-to-Prompt [18] to generate large-scale image-instruction pairs. To reduce data dependency, some work [5, 8] extracts consecutive video frames as editing samples to simulate real-world editing dynamics. Recent studie  [33] adapt the InstructPix2Pix framework to autoregressive models through fine-tuning on paired datasets.
@@ 3.1 Preliminary
$P(z|c)=\prod_{i=1}^{N}P(z_{i}|z_{<i},c)\vspace{-1mm}$ (1)
where $N=h\times w$. During generation, the concatenated sequence $[c;{z}_{<i}]$ is processed through stacked transformer layers with causal masking to enforce autoregressive constraints. At each layer $l$, the self-attention mechanism is:
$\mathbf{A}_{l}=\text{Softmax}\left(\frac{\mathbf{Q}_{l}\mathbf{K}_{l}^{\top}}{\sqrt{d_{k}}}\right)\vspace{-2mm}$ (2)
@@ 3.3 Implicit Structure Locking (ISLock)
$z_{i}^{\text{edit}}=\arg\min_{k\in\mathcal{W}_{i}}s^{(k)}\vspace{-1mm}$ (5)
@@
Method T2I Model Distance $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ Whole $\uparrow$ Edited $\uparrow$
@@ 4.1 Experimental Setups
Benchmarks. We build our method ISLock on LlamaGen [46] to generate images at a resolution of 512$\times$512. By default, we set $K=150,\tau=1.0$ as our hyperparameters. To establish a rigorous evaluation framework for training-free AR image editing, we designed a generation-to-editing pipeline that aligns with our method’s operational paradigm. Given the intrinsic requirement of our approach to process AR-generated images, we evaluated our method on a curated subset of the PIE-Bench dataset [23], the standard benchmark for image editing. While existing methods typically support all 10 editing types in PIE-Bench, we focused on 5 fundamental categories compatible with our AR training-free paradigm: object replacement, object addition, object removal, style transfer, and attribute modification. Each case in the curated subset of the PIE-Bench [22] includes paired original and editing prompts, we first used LlamaGen to generate images based on original prompts, and then applied ISLock to edit these images according to editing prompts.
@@ 4.2 Experimental Results
We also implement two simple baselines: Naive Prompt Modification (NPM), which directly modifies target words in the prompt, and PnP-AR [18], which replaces attention maps generated from the original image during editing. As shown in Table 1, these baselines struggle to balance structural consistency with background preservation, exhibiting significantly higher Structure Distance scores (113.95 and 103.94, respectively) and inferior background fidelity. By contrast, ISLock achieves satisfactory structural distance (31.79$\downarrow$, second only to inversion-based methods), indicating superior strctural preservation during editing operations. Furthermore, our method maintains favorable background fidelity, performing comparably to leading inversion-free instruction-based editing methods (InstructPix2Pix [2], MGIE [12]) in PSNR and SSIM metrics. Notably, ISLock attains high CLIP similarity scores for both whole image and edited regions (24.19&21.33, only second to PnPInversion and P2P), highlighting its ability to align edited content with target text prompts while preserving semantic consistency.
@@ 4.3 Ablation Study
Effect of Window Size $|\mathcal{W}|$. As shown in Fig. 7 and Table 2, the ablation study on window size $|\mathcal{W}|$ reveals its critical role in balancing structural preservation and model generation flexibility. When $|\mathcal{W}|=1$, our method degenerates into naïve prompt modification, resulting in less controlled edits. As $|\mathcal{W}|$ increases, the structural distance between the edited image and the original decreases, indicating better preservation of image composition. However, this comes at the cost of reduced generative flexibility. As seen in Fig. 7, when $|\mathcal{W}|$ is too large, a “happy dog” with its tongue out is modified to have a closed mouth, and the chocolate coating on a cake is replaced with a plain bread layer. This effect is further reflected in the drop in CLIP Similarity (Table 2), suggesting weaker alignment with the target prompt. In contrast, our dynamic window strategy, shown in the last column of Fig. 7, achieves a better balance between structural consistency and semantic alignment, resulting in the best Structure Distance / CLIP Similarity ratio.
@@ 5 Conclusion
In this work, we address the fundamental challenge of text-guided image editing in autoregressive (AR) models without modifying model parameters or relying on explicit attention manipulation. By introducing Implicit Structure Locking (ISLock), we enable training-free image editing through a novel candidate matching protocol named Anchor Token Matching (ATM), which aligns the structure of the original image while preserving generative autonomy. Unlike direct attention map transplantation, our method ISLock ensures structure preservation as a natural consequence of latent space structure, allowing AR models to maintain spatial coherence for various editing tasks. Extensive experiments demonstrate that ISLock achieves competitive performance in structure-aware autoregressive editing, bridging the gap between AR models and diffusion models.
@@ B.3 Attention Locality
We observe that in autoregressive models such as LlamaGen, tokens tend to allocate higher attention weights to those adjacent to their positions during attention computation. As shown in Figure 11, the attention score assigned by the current token decreases as the distance from the current token increases. However, the attention score periodically increases at intervals of 32 tokens. This phenomenon occurs because, when generating 512$\times$512 images, LlamaGen employs a VQ-Autoencoder to encode the image into a 32$\times$32 latent space. Tokens located at multiples of 32 positions away from the current token reside in the same column in the latent space, resulting in higher attention scores at these intervals. This also explains why the cross-attention from image tokens to text tokens in Figure 4 of main paper shows that the earliest image tokens have the highest attention scores.

### v2

@@ 2 Related Work
Text-guided Image Editing. Text-guided image editing aims to modify image content based on semantic prompts while preserving irrelevant regions. Traditional GAN-based methods [16, 30, 36, 36] achieve local editing by optimizing in the GAN latent space with CLIP guidance [22, 37], but their performance is constrained by the capacity of pre-trained GANs. Recent diffusion-based approaches have become mainstream, yet they often require complex inversion and latent optimization processes to balance fidelity and controllability. For example, optimization-based inversion methods [31, 55, 26, 17] refine latent noise for accurate reconstruction and manipulate cross-attention maps (e.g., Prompt-to-Prompt [18]) to preserve structure. InstructPix2Pix [2] bypasses inversion by training on paired edit data, but relies on Prompt-to-Prompt [18] to generate large-scale image-instruction pairs. To reduce data dependency, some work [5, 8] extracts consecutive video frames as editing samples to simulate real-world editing dynamics. Recent studie  [34] adapt the InstructPix2Pix framework to autoregressive models through fine-tuning on paired datasets.
@@ 3.1 Preliminary
$P(z|c)=\prod_{i=1}^{N}P(z_{i}|z_{<i},c)\vskip-2.84526pt$ (1)
where $N=h\times w$. During generation, the concatenated sequence $c;{z}_{<i}]$ is processed through stacked transformer layers with causal masking to enforce autoregressive constraints. At each layer $l$, the self-attention mechanism is:
$\mathbf{A}_{l}=\text{Softmax}\left(\frac{\mathbf{Q}_{l}\mathbf{K}_{l}^{\top}}{\sqrt{d_{k}}}\right)\vskip-5.69054pt$ (2)
@@ 3.3 Implicit Structure Locking (ISLock)
$z_{i}^{\text{edit}}=\arg\min_{k\in\mathcal{W}_{i}}s^{(k)}\vskip-2.84526pt$ (5)
@@
Method
T2I
Model
Distance $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ Whole $\uparrow$ Edited $\uparrow$
@@ 4.1 Experimental Setups
Benchmarks. We build our method ISLock on LlamaGen [48] to generate images at a resolution of $\times512$. By default, we set $K=150,\tau=1.0$ as our hyperparameters. To establish a rigorous evaluation framework for training-free AR image editing, we designed a generation-to-editing pipeline that aligns with our method’s operational paradigm. Given the intrinsic requirement of our approach to process AR-generated images, we evaluated our method on a curated subset of the PIE-Bench dataset [24], the standard benchmark for image editing. While existing methods typically support all 10 editing types in PIE-Bench, we focused on 5 fundamental categories compatible with our AR training-free paradigm: object replacement, object addition, object removal, style transfer, and attribute modification. Each case in the curated subset of the PIE-Bench [23] includes paired original and editing prompts, we first used LlamaGen to generate images based on original prompts, and then applied ISLock to edit these images according to editing prompts.
@@ 4.2 Experimental Results
[…]
@@ 4.3 Ablation Study
Effect of Window Size $|\mathcal{W}|$. As shown in Fig. 7 and Table 2, the ablation on window size $|\mathcal{W}|$ reveals a critical trade-off between structural preservation and generative flexibility. A larger $|\mathcal{W}|$ enhances structural preservation (lower structural distance) but can excessively constrain the model, thereby reducing generative flexibility and alignment with the target prompt (lower CLIP Similarity). For instance, an overly large window might incorrectly alter a dog’s expression or a cake’s topping, as seen in Fig. 7. In contrast, our proposed dynamic window strategy achieves a superior balance, resulting in the best Structure Distance / CLIP Similarity ratio.
@@ 5 Conclusion
In this work, we address the fundamental challenge of text-guided image editing in autoregressive (AR) models without modifying model parameters or relying on explicit attention manipulation. By introducing Implicit Structure Locking (ISLock), we enable training-free image editing through a novel candidate matching protocol named Anchor Token Matching (ATM), which aligns the structure of the original image while preserving generative autonomy. Our method ISLock ensures structure preservation as a natural consequence of latent space structure, allowing AR models to maintain spatial coherence for various editing tasks. Extensive experiments demonstrate that ISLock achieves competitive performance in structure-aware autoregressive editing, bridging the gap between AR models and diffusion models.
Acknowledgements
This work was supported by NSFC (NO. 62225604) and Youth Foundation (62202243).We acknowledge “Science and Technology Yongjiang 2035” key technology breakthrough plan project (2024Z120). We also acknowledge the support of the Shenzhen Science and Technology Program (JCYJ20240813114237048), the project PID2022-143257NB-I00 funded by the Spanish Government through MCIN/AEI/10.13039/501100011033 and FEDER, and the Generalitat de Catalunya CERCA Program. Computation is supported by the Supercomputing Center of Nankai University (NKSC).
@@ B.3 Attention Locality
We observe that in autoregressive models such as LlamaGen, tokens tend to allocate higher attention weights to those adjacent to their positions during attention computation. As shown in Figure 11, the attention score assigned by the current token decreases as the distance from the current token increases. However, the attention score periodically increases at intervals of 32 tokens. This phenomenon occurs because, when generating $\times512$ images, LlamaGen employs a VQ-Autoencoder to encode the image into a $\times32$ latent space. Tokens located at multiples of 32 positions away from the current token reside in the same column in the latent space, resulting in higher attention scores at these intervals. This also explains why the cross-attention from image tokens to text tokens in Figure 4 of main paper shows that the earliest image tokens have the highest attention scores.